# VITA: Vision-to-Action Flow Matching Policy

### v2

@@ Abstract
Conventional flow matching and diffusion-based policies sample through iterative denoising from standard noise distributions (e.g., Gaussian), and require conditioning mechanisms to incorporate visual information during the generative process, incurring substantial time and memory overhead. To reduce the complexity, we develop VITA (VIsion-To-Action policy), a noise-free and conditioning-free policy learning framework that directly maps visual representations to latent actions using flow matching. VITA treats latent visual representations as the source of the flow, thus eliminating the need of conditioning. As expected, bridging vision and action is challenging, because actions are lower-dimensional, less structured, and sparser than visual representations; moreover, flow matching requires the source and target to have the same dimensionality. To overcome this, we introduce an action autoencoder that maps raw actions into a structured latent space aligned with visual latents, trained jointly with flow matching. To further prevent latent space collapse, we propose flow latent decoding, which anchors the latent generation process by backpropagating the action reconstruction loss through the flow matching ODE (ordinary differential equations) solving steps. We evaluate VITA on 8 simulation and 2 real-world tasks from ALOHA and Robomimic. VITA outperforms or matches state-of-the-art generative policies, while achieving $1.5{\times}$-$2.3{\times}$ faster inference compared to conventional methods with conditioning. Project page: VITA.
@@ 1 Introduction
Figure 1: A comparison between VITA and conventional flow matching and diffusion policies. Unlike conventional methods that sample noise from standard distributions and inject source input modalities via conditioning during generation, VITA poses no constraints on the source distributions, and flow directly from latent visual representations to latent actions, eliminating the need for conditioning.
@@
Flow matching and diffusion models have demonstrated remarkable success across a wide range of cross-modal generation tasks, from text-to-image generation stablediffusion ; dit ; sit ; crossflow ; flowtok ; conddiffusion , text-to-video generation video-diff ; li2023videogen ; jin2024pyramidal , to visuomotor (vision-to-action) policy chi2023diffusion ; ren2024diffusion ; gao2025ril ; su2025dense ; fmp-3d-consist ; fmp-multisupport ; fmp-riemannian ; black2410pi0 . Conventional flow matching and diffusion methods  lipman2023flowmatchinggenerativemodeling ; sohl2015deep generate samples by starting with noise sampled from a basic source distribution (often Gaussian) and progressively “denoising” them into the target modality, requiring additional conditioning mechanisms  stablediffusion ; conddiffusion ; chi2023diffusion ; dit-policy to inject visual modality at denoising steps, which incurs increased time and space complexity crossflow ; flowtok .
@@
Clearly, minimizing complexity is critical for effective real-time robot control. To overcome the inefficiencies inherent to conditioning mechanisms in conventional methods, we develop VITA (VIsion-To-Action flow matching), a noise-free policy learning framework that directly maps visual representations to latent actions. As depicted in  Figure˜1, unlike conventional methods that flow from a Gaussian prior and inject visual contexts via conditioning mechanisms during generation, VITA poses no constraints on the source distribution, and directly flows from latent visual representations, obviating the need for injecting visual inputs via conditioning modules. Consequently, VITA significantly reduces inference latency and memory overhead, simplifying the network architecture.
@@
Learning vision-to-action flow matching, however, presents several new challenges. In general, bridging two distinct modalities is inherently difficult crossflow ; flowtok , particularly in robot learning where action data is relatively limited, unstructured and sparse, whereas the visual representations exhibit rich structures and semantics and have far higher dimensions. Additionally, flow matching requires that the source and target have equal dimensionalities, which prevents directly pairing raw actions with visual representations.
@@
To address these challenges, we propose two key designs for VITA. 1) Learned target latent actions for flow matching. We introduce a structured latent action space, learned via an action autoencoder, that ‘lifts’ action representations to match the higher dimensionality of visual representations and serves as a structured target distribution for flow matching. The action encoder up-samples raw actions into target latent actions, and a decoder reconstructs raw actions from these latents. 2) Flow latent decoding. In conventional flow matching methods such as latent diffusion for image generation (stablediffusion, ), the target latent space can be pre-trained with abundant image data and then frozen as reliable flow matching targets; in contrast, we show that a pre-trained and frozen latent action space for learning vision-to-action flows yields poor performance, since action data can be too sparse and limited to learn reliable target latents and frozen latent spaces cannot be corrected. It is therefore plausible to jointly train the flow model with the action autoencoder. To prevent model collapse of target latent action space (which happens by naively reducing the flow matching and autoencoder loss), we introduce flow latent decoding which backpropagates the reconstruction loss of latent actions generated by solving flow matching ordinary differential equations (ODE), anchoring the latent generation process using ground-truth actions. This approach also bridges the training-inference gap of latent actions: during training, the flow matching model learns to match the targets given by the action encoder, and the decoder is trained to reconstruct these targets. whereas at test time, the decoder must reconstruct ODE-generated latent actions.
We evaluate VITA on both real-world and simulated tasks using ALOHA chuang2024active ; fu2024mobile ; zhao2024aloha and Robomimic robomimic2021 . Our experiments demonstrate that VITA outperforms or matches state-of-the-art generative policies. We implement VITA with only MLP layers. Compared to state-of-the-art flow matching and diffusion policies that predominantly leverage complex architecture such as transformers sit or U-Net unet , and introduce conditioning modules zhao2023learning ; fmp-3d-consist , the MLP-only and noise-free VITA policy achieves, $1.5{\times}$-$2.3{\times}$ faster inference compared to conventional flow matching. To our knowledge, VITA is the first MLP-only flow matching policy to succeed on tasks as challenging as ALOHA bi-manual manipulation.
@@
Noise-Free Flow Matching for Visuomotor Learning. We propose VITA, a noise-free policy that directly evolves latent visual representations into latent actions via flow matching. To bridge the modality gap, VITA incorporates an action autoencoder that maps raw actions into structured latent spaces aligned with visual representations. To prevent latent space collapse, we propose flow latent decoding, which anchors latent generation using ground-truth actions by backpropagating the action reconstruction loss through the flow matching ODE (ordinary differential equations) solving steps.
An Efficient Policy Architecture. By treating vision as the source distribution, VITA obviates costly conditioning mechanisms required by conventional diffusion and flow matching policies. VITA also enables a lightweight MLP-only implementation that achieves superior performance on challenging bi-manual manipulation tasks with high-dimensional visual inputs.
State-of-the-Art Performance in Real and Simulated Environments. We validate VITA’s effectiveness on 8 simulation and 2 real-world tasks. VITA surpasses or matches state-of-the-art generative policies in success rates while delivering $1.5{\times}$-$2.3{\times}$ faster inference compared to conventional flow matching approaches.
@@ 2 Related Work
Imitation Learning for Visuomotor Policy. Imitation learning enables robots to learn complex behaviors by mimicking expert demonstrations. Behavior cloning is a prominent imitation learning paradigm that learns a policy that maps observations to actions (zhao2023learning, ; lee2024interact, ) via supervised learning. Recent advancements in behavioral cloning have widely adopted autoregressive modeling (fu2024context, ; carp, ; su2025dense, ) and generative modeling (chi2023diffusion, ). Generative modeling learns a conditional distribution of actions given an observation, leveraging conditional variational autoencoders (CVAEs) (zhao2023learning, ; lee2024interact, ), diffusion (dit-policy, ; chi2023diffusion, ), or flow matching (fmp-affordance, ; fmp-3d-consist, ). Generative models ubiquitously require conditioning modules (e.g., cross-attention (dit-policy, ), AdaLN (dit-policy, ), FiLM (perez2018film, ; chi2023diffusion, )) to inject observations at each step of the generation process. VITA removes the visual conditioning module by developing a noise-free and conditioning-free vision-to-action flow.
Diffusion and Flow matching for Generative Modeling. Diffusion and flow matching models have shown strong performance across diverse generative tasks, such as image and video generation (stablediffusion, ; video-diff, ), and visuomotor policies (chi2023diffusion, ). Unlike diffusion which samples from a Gaussian distributions, flow matching theoretically places no constraints on the choice of source distribution (tong2024bridgesscore, ). A few works have explored this property to learn the direct transport within the same modality (i2i-flow-interpolants, ; i2i-flow-sim-free, ). Recently,  crossflow and  flowtok extended this to more challenging cross-modal generation between text and image. VITA focuses on learning to bridge vision and action for visuomotor control, where the action modality presents unique challenges because of limited data and its unstructured nature. Different from flow matching for image generation, which typically pre-trains and freezes the latent image space when learning the flow (stablediffusion, ; flowtok, ; crossflow, ), VITA resorts to a fully end-to-end pipeline training to effectively learn the latent action space from limited and sparse action data along with flow matching. We propose flow latent decoding to backpropagate action reconstruction losses through the the latent action generation process (ODE solving steps) during training.
@@ 3 Preliminaries
Flow matching models learn to transport samples from a source probability distribution $p_{0}$ to a target distribution $p_{1}$ by learning a velocity vector field $v_{\theta}$ (lipman2023flowmatchinggenerativemodeling, ; rectified-flow, ). This generative process is defined by an ordinary differential equation (ODE), $\frac{d\bm{z}_{t}}{dt}=v_{\theta}(\bm{z}_{t},t)$, where $t\in[0,1]$ is a continuous time variable, $\bm{z}_{t}$ represents a sample at time $t$, and the model $v_{\theta}$ is a neural network. The goal of training is for $v_{\theta}$ to generate a path that transforms a sample $\bm{z}_{0}\sim p_{0}$ into a corresponding sample $\bm{z}_{1}\sim p_{1}$.
For a simple, straight probability path between two samples, the interpolated state at time $t$ is defined as $\bm{z}_{t}=(1-t)\bm{z}_{0}+t\bm{z}_{1}$ The corresponding ground-truth velocity of this path given by $u_{t}=\frac{d\bm{z}_{t}}{dt}=\bm{z}_{1}-\bm{z}_{0}$. The network $v_{\theta}$ is trained by minimizing the Mean Squared Error (MSE) between its predicted velocity and the ground-truth velocity:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t,\bm{z}_{0},\bm{z}_{1}}\left[\left\|v_{\theta}(\bm{z}_{t},t)-(\bm{z}_{1}-\bm{z}_{0})\right\|^{2}\right]$ (1)
During inference, starting from a source sample $\bm{z}_{0}$, the model generates a target sample $\hat{\bm{z}}_{1}$ by solving the ODE initial value problem from $t=0$ to $t=1$:
$\hat{\bm{z}}_{1}=\bm{z}_{0}+\int_{0}^{1}v_{\theta}(\bm{z}_{t},t)\,dt$ (2)
Notably, conventional flow matching or diffusion policies for visuomotor control generate actions by evolving a sample from a simple prior distribution, such as Gaussian noise ($\bm{z}_{0}\sim\mathcal{N}(0,I)$), and require conditioning on visual inputs to guide the generation process (fmp-affordance, ; chi2023diffusion, ; dit-policy, ). However, flow matching does not assume the source distribution must be simple noise; it can theoretically connect two arbitrary distributions. We exploit this flexibility to re-formulate generative visuomotor policy learning: we directly use the latent distribution of visual observations as the source distribution $p_{0}$, yielding a noise-free flow matching process that eliminates the need for costly conditioning modules and therefore boosts time and space efficiency.
@@ 4.1 Flowing from Vision to Action
VITA learns a policy $\pi(A|O)$ that directly maps observations $O$ to a corresponding sequence of future actions $A$. The observations $O$ encompasses raw visual inputs $I\in\mathbb{R}^{H\times W\times C}$ and, optionally, the robot’s proprioceptive state vector $S$. Actions are represented as temporal sequences over a prediction horizon, formally defined as $A\in\mathbb{R}^{T_{\text{pred}}\times D_{\text{action}}}$, where $T_{\text{pred}}$ is the prediction horizon and $D_{\text{action}}$ is the dimensionality of the action space. We employ action chunking with $T_{\text{pred}}>1$ to enhance temporal consistency  (zhao2023learning, ).
Unlike conventional flow matching policies that generate actions through conditional denoising of random noise given visual inputs, VITA formulates the visuomotor policy as an unconditional flow between the source distribution of latent visual representations $p_{0}$ and the target distribution of latent actions $p_{1}$. The flow matching framework learns to evolve a latent visual representation $\bm{z}_{0}$ into an latent action representation $\bm{z}_{1}$ by learning a velocity field $v_{\theta}(\bm{z}_{t},t)$, where $t\in[0,1]$ represents the continuous time parameter interpolating between the source state at $t=0$ and the target state at $t=1$. Critically, flow matching requires $\bm{z}_{0}$ and $\bm{z}_{1}$ to share identical dimensionality, necessitating the construction of a latent action space that matches the dimensionality of visual representations (Section˜4.3).
During inference, the current observation $O_{\text{curr}}$ is first encoded into its latent visual representation $\bm{z}_{0}=\mathcal{E}_{v}(O_{\text{curr}})$. This latent vector is subsequently evolved into a predicted action latent representation, $\hat{\bm{z}}_{1}$, by numerically solving the ODE from $t=0$ to $t=1$ using the learned velocity field $v_{\theta}$: $\hat{\bm{z}}_{1}=\bm{z}_{0}+\int_{0}^{1}v_{\theta}(\bm{z}_{t},t)\,dt$. In other words, $\hat{\bm{z}}_{1}$ is obtained by numerically solving the flow matching ODE, and serves as an approximation to the target latent $\bm{z}_{1}$. The resulting latent action vector $\hat{\bm{z}}_{1}$ is then decoded through the action decoder to yield the final action sequence $\hat{A}=\mathcal{D}_{a}(\hat{\bm{z}}_{1})$.

### v3

@@ Abstract
Conventional flow matching and diffusion-based policies sample through iterative denoising from standard noise distributions (e.g., Gaussian), and require conditioning modules to repeatedly incorporate visual information during the generative process, incurring substantial time and memory overhead. To reduce the complexity, we develop VITA (VIsion-To-Action policy), a noise-free and conditioning-free flow matching policy learning framework that directly flows from visual representations to latent actions. Since the source of the flow is visually grounded, VITA eliminates the need of visual conditioning during generation. As expected, bridging vision and action is challenging, because actions are lower-dimensional, less structured, and sparser than visual representations; moreover, flow matching requires the source and target to have the same dimensionality. To overcome this, we introduce an action autoencoder that maps raw actions into a structured latent space aligned with visual latents, trained jointly with flow matching. To further prevent latent space collapse, we propose flow latent decoding, which anchors the latent generation process by backpropagating the action reconstruction loss through the flow matching ODE (ordinary differential equation) solving steps. We evaluate VITA on 9 simulation and 5 real-world tasks from ALOHA and Robomimic. VITA achieves $1.5{\times}$-$2{\times}$ faster inference compared to conventional methods with conditioning modules, while outperforming or matching state-of-the-art policies. Codes, datasets, and demos are available at our project page: VITA.
@@ 1 Introduction
Figure 1: A comparison between VITA and conventional flow matching and diffusion policies. Unlike conventional methods that sample noise from standard distributions and inject input modalities via conditioning, VITA poses no constraints on the source distribution, and flows directly from latent visual representations to latent actions, eliminating the need for conditioning modules.
@@
Flow matching and diffusion models have demonstrated remarkable success across a wide range of cross-modal generation tasks, from text-to-image generation (stablediffusion; dit; sit; crossflow; flowtok; conddiffusion), text-to-video generation (video-diff; li2023videogen; jin2024pyramidal), to visuomotor (vision-to-action) policies (chi2023diffusion; ren2024diffusion; gao2025ril; su2025dense; fmp-3d-consist; fmp-multisupport; fmp-riemannian; black2410pi0). Conventional flow matching and diffusion methods  (lipman2023flowmatchinggenerativemodeling; sohl2015deep) generate samples by starting with noise sampled from a basic source distribution (often Gaussian) and progressively “denoising” them into the target modality. This process requires repeatedly injecting visual information at each denoising step through additional conditioning modules (stablediffusion; conddiffusion; chi2023diffusion; dit-policy), resulting in substantial time and space overheads  (crossflow; flowtok). In particular, cross-attention, AdaLN (dit), or FiLM (perez2018film) are often used for visual conditioning. Cross-attention incurs quadratic time and space complexity, while AdaLN and FiLM avoid quadratic complexity but require extra modulation networks to generate feature-wise parameters at every denoising step.
@@
Minimizing complexity is essential for real-time robot control, e.g., Pi-0.5 (pi05) operates at  50 Hz and Helix (helix-figure) at up to  200 Hz, imposing stringent requirements on inference latency. The primary objective of this paper is to overcome the inefficiencies inherent to conditioning mechanisms in conventional flow matching methods. To this end, we develop VITA (VIsion-To-Action policy), a noise-free flow matching policy learning framework that directly maps visual representations to latent actions. As depicted in  Figure˜1, unlike conventional methods that flow from a Gaussian prior, VITA imposes no constraints on the source distribution and flows directly from visual latents, obviating the need for repeated visual conditioning during the flow. Consequently, VITA significantly reduces time and space overheads, and simplifies network architectures.
@@
Learning vision-to-action flow matching, however, presents several new challenges. Bridging two distinct modalities is inherently difficult (crossflow), particularly in robotics where action data is limited, unstructured, and sparse, whereas visual representations exhibit rich structures and semantics, and far higher dimensions. Additionally, flow matching requires that the source and target have equal dimensionalities, which prevents directly pairing raw actions with visual representations.
@@
To address these challenges, we propose two key designs for VITA. 1) Learned target latent actions for flow matching. We introduce a structured latent action space, learned via an action autoencoder, that ‘lifts’ action representations to match the higher dimensionality of visual representations and serves as a structured target distribution for flow matching. The action encoder up-samples raw actions into target latent actions, and a decoder reconstructs raw actions from these latents. 2) Flow latent decoding. In conventional flow matching methods such as latent diffusion for image generation (stablediffusion), the target latent space can be pre-trained with abundant image data and then frozen as reliable flow matching targets; in contrast, we show that a pre-trained and frozen latent action space for learning vision-to-action flows yields poor performance (discussed in Appendix B.2), since action data is too sparse and limited to learn reliable targets and frozen targets cannot be corrected. It is therefore plausible to jointly train the flow model with the action autoencoder. To prevent model collapse of the target latent action space (which happens by naively reducing flow matching and autoencoder losses), we introduce flow latent decoding which backpropagates the reconstruction loss of latent actions generated by solving flow matching ordinary differential equations (ODEs), anchoring latent generation using ground-truth actions. This approach also bridges the training-inference gap of latent actions: during training, the flow matching model learns to match the targets given by the action encoder, and the decoder is trained to reconstruct these targets, whereas at test time, the decoder must reconstruct ODE-generated latent actions.
We evaluate VITA on both real-world and simulated tasks using ALOHA (chuang2024active; fu2024mobile; zhao2024aloha) and Robomimic (robomimic2021). VITA achieves $1.5{\times}-2{\times}$ faster inference and 18.6%-28.7% lower memory usage compared to conventional flow matching with similar model sizes, while outperforming or matching state-of-the-art policies in success rates. Additionally, compared to state-of-the-art methods that necessitate complex architectures (e.g., transformers (sit)), VITA naturally simplifies architecture designs. For instance, with vector-based visual representations, VITA reduces the flow-matching network to a conditioning-free vector-to-vector mapping, allowing for the use of simple MLPs; with higher-dimensional grid-based visual representations, VITA scales to more complex architectures such as transformers while eliminating costly conditioning modules like cross-attention.
@@
Noise-Free Flow Matching for Visuomotor Learning. We propose VITA, a noise-free policy that directly evolves latent visual representations into latent actions via flow matching. VITA learns a structured latent action space aligned with visual representations to bridge the modality gap. To prevent latent collapse during end-to-end training, we propose flow latent decoding, which refines latent actions by backpropagating through the flow matching ODE solving steps.
Efficient Policy Architectures. By visually grounding the source of the flow, VITA obviates costly conditioning required by flow matching policies to repeatedly inject visual inputs. VITA enables lightweight implementations. To our knowledge, VITA is the first MLP-only flow matching policy to succeed on tasks as challenging as ALOHA bimanual manipulation.
State-of-the-Art Efficiency and Performance. We validate VITA’s efficiency and performance on 9 simulation and 5 real-world tasks spanning both bimanual and single-arm manipulation. VITA delivers $1.5{\times}$-$2{\times}$ faster inference and 18.6%-28.7% lower memory usage compared to conventional flow matching, while surpassing or matching state-of-the-art policies in success rates.
@@ 2 Related Work
Imitation Learning for Visuomotor Policy. Imitation learning enables robots to learn complex behaviors by mimicking expert demonstrations. Behavior cloning is a prominent imitation learning paradigm that learns a policy that maps observations to actions (zhao2023learning; lee2024interact) via supervised learning. Recent advancements in behavioral cloning have widely adopted autoregressive modeling (fu2024context; carp; su2025dense) and generative modeling (chi2023diffusion). Generative modeling learns a conditional distribution of actions given an observation, leveraging conditional variational autoencoders (CVAEs) (zhao2023learning; lee2024interact), diffusion (dit-policy; chi2023diffusion), or flow matching (fmp-affordance; fmp-3d-consist). Generative models ubiquitously require conditioning modules (e.g., cross-attention (dit-policy), AdaLN (dit-policy), FiLM (perez2018film; chi2023diffusion)) to inject observations at each step of the generation process. VITA removes the visual conditioning module by developing a noise-free vision-to-action flow.
Diffusion and Flow Matching for Generative Modeling. Unlike diffusion which samples from a Gaussian distributions, flow matching theoretically places no constraints on the choice of source distribution (tong2024bridgesscore). A few works have explored this property to learn the direct transport within the same modality (i2i-flow-interpolants; i2i-flow-sim-free). Recently,  crossflow and  flowtok extended this to more challenging cross-modal generation between text and image. VITA learns to bridge vision and action for visuomotor control, where the action modality presents unique challenges because of limited data and its unstructured nature. Different from flow matching for image generation, which typically pre-trains and freezes the latent image space when learning the flow (stablediffusion; crossflow), VITA resorts to a fully end-to-end pipeline training to effectively learn the latent action space from limited and sparse action data along with flow matching. We propose flow latent decoding to backpropagate action reconstruction losses through the latent action generation process (ODE solving steps) during training.
@@ 3 Preliminaries
Flow matching models learn to transport samples from a source distribution $p_{0}$ to a target distribution $p_{1}$ by learning a velocity vector field $v_{\theta}$ (rectified-flow). The generative process is defined by an ODE $\frac{d\mathbf{z}_{t}}{dt}=v_{\theta}(\mathbf{z}_{t},t)$  (lipman2023flowmatchinggenerativemodeling; anypose-ode), where $t\in[0,1]$ is continuous time, and $\mathbf{z}_{t}$ denotes a sample at time $t$. The goal is for the learned flow to transport $\mathbf{z}_{0}\sim p_{0}$ to $\mathbf{z}_{1}\sim p_{1}$.
Training. For a linear interpolation between two samples, the interpolation path is $\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\mathbf{z}_{1}$. The ground-truth velocity along this path is $\frac{d\mathbf{z}_{t}}{dt}=\mathbf{z}_{1}-\mathbf{z}_{0}$. The flow matching loss trains $v_{\theta}$ to match this supervised vector field:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t,\,\mathbf{z}_{0},\,\mathbf{z}_{1}}\left[\left\|v_{\theta}(\mathbf{z}_{t},t)-(\mathbf{z}_{1}-\mathbf{z}_{0})\right\|^{2}\right].$ (1)
Inference. Given a source sample $\mathbf{z}_{0}$, a target sample is obtained by solving the ODE from $t=0$ to $t=1$: $\hat{\mathbf{z}}_{1}=\mathbf{z}_{0}+\int_{0}^{1}v_{\theta}(\mathbf{z}_{t},t)\,dt$. In practice, we apply an Euler solver with $K$ discretization steps, yielding updates of the form $\mathbf{z}_{t_{k+1}}=\mathbf{z}_{t_{k}}+\Delta t\,v_{\theta}(\mathbf{z}_{t_{k}},t_{k})$, where $\Delta t=1/K$.
@@ 4.1 Flowing from Vision to Action
VITA learns a policy $\pi(A|O)$ that directly maps observations $O$ to a corresponding sequence of future actions $A$. The observations $O$ encompasses raw visual inputs $I\in\mathbb{R}^{H\times W\times C}$ and, optionally, the robot’s proprioceptive states $S$. Actions are represented as temporal sequences over a prediction horizon, formally defined as $A\in\mathbb{R}^{T_{\text{pred}}\times D_{\text{action}}}$, where $T_{\text{pred}}$ is the prediction horizon and $D_{\text{action}}$ is the dimensionality of the action space. We employ action chunking with $T_{\text{pred}}>1$ to enhance temporal consistency  (zhao2023learning).
Conventional vs. VITA Flow Matching. Conventional flow matching policies generate actions by evolving samples from a noise prior, typically $\mathbf{z}_{0}\sim\mathcal{N}(0,I)$. To incorporate visual information, these models learn a conditional velocity field $v_{\theta}(\mathbf{z}_{t},t\mid O)$, requiring conditioning modules (e.g., cross-attention) to inject observations $O$ at every denoising step. In contrast, VITA directly treats the visual latent as the source of the flow $\mathbf{z}_{0}$. Because the flow is visually grounded at the source, VITA learns a conditioning-free velocity field for flow matching, $v_{\theta}(\mathbf{z}_{t},t)$, eliminating the need for repetitive conditioning and yielding a noise-free framework with enhanced efficiency.
Critically, flow matching requires $\bm{z}_{0}$ and $\bm{z}_{1}$ to share identical dimensionality, necessitating the construction of a latent action space that matches the dimensionality of visual representations (Section˜4.3). During inference, the current observation $O_{\text{curr}}$ is first encoded into its latent visual representation $\bm{z}_{0}=\mathcal{E}_{v}(O_{\text{curr}})$, which is subsequently evolved into a predicted latent action representation, $\hat{\bm{z}}_{1}$, by numerically solving the ODE from $t=0$ to $t=1$ using the learned velocity field $v_{\theta}$. In other words, $\hat{\bm{z}}_{1}$ is an approximation to the target latent $\bm{z}_{1}$. The resulting latent action $\hat{\bm{z}}_{1}$ is then decoded through the action decoder to yield the final action sequence $\hat{A}=\mathcal{D}_{a}(\hat{\bm{z}}_{1})$.